OpenStudy (cicilybailey14):

I medal... Suppose y varies inversely with x and x = 12 when y = 3. What is the equation of the inverse variation? y/x = 36 12 = 3/y y = 12/3 y = 36/x

OpenStudy (danjs):

Y varies inversely as x means. y = k/x where k is the constant of variation. Use the x and y value they gave you to figure out what the k value is

OpenStudy (danjs):

\[3 = \frac{ k }{ 12 }\]

OpenStudy (danjs):

Then use that k value in: \[y = \frac{ k }{ x }\] for your final answer

OpenStudy (danjs):

\[y = \frac{ 12 * 3 }{ x }\]

OpenStudy (cicilybailey14):

@DanJS so C?